PHYS10471 2. The following equation occurs in conditional probability: a) 5 marks] B marks 2 marks] b) An electronic system contains in components which are connected in series and they i) Describe the meaning of each term in the equation ii) Write down an expression for the unconditional probability P() in terms of iii) Describe the implications omrix)>MYIX). quantities in the above equation. function independently of each other. The length of time for each component until failure follows an exponential distribution density function with mean Determine the mean and variance of the length of time until failure of the system. [5 marks] c) Suppose that 10% of people in a certain population have the eye disease glaucoma. For persons who have glaucoma, measurements of eye pressure x will be normally distributed with a mean of 25 mm-Hg and a variance of 1 mm-Hg. For persons who do not have glaucoma, the pressure will be normally distributed with a mean of 20 mm-Hg and a variance of 1 mm-Hg. Suppose that a person is selected from the population and her eye pressure measured. Determine the conditional probability that the person has glaucoma given [5 marks] i) pressure x. ii) For what value of x is the conditional probability in i) greater than 1/2? 5 marks]
Answer #1

2 a)

(i)

P(X|Y) = conditional probability that event X occurs given that event Y occurs

P(Y|X) = conditional probability that event Y occurs given that event X occurs

P(X) = unconditional probability that event X occurs

P(Y|overline{X}) = conditional probability that event Y occurs given that event X does not occur

P(overline{X}) = unconditional probability that event X does not occur

(ii)

P(Y) = P(YcapX) + P(Ycapoverline{X})

Rightarrow P(Y) = P(Y|X)P(X) + P(Y|overline{X})P(overline{X})

P(Y) = P(X)}

(iii)

P(Y|overline{X}) = P(Y|X)

Rightarrow conditional probability that event Y occurs given that event X does not occur

= conditional probability that event Y occurs given that event X occurs

Now ; P(Y|overline{X}) = P(Y|X)

herefore P(X|Y)=rac{P(Y|X)P(X)}{P(Y|X)P(X)+P(Y|overline{X})P(overline{X})}

P(X) P(X) + P(X) P(X|Y) = P(YX)PX) + P(Y|X)P(X)

i.e ; X and Y are independent events   

2 b)

2> -- X, denete the lensths ef bime until foilure fon Ist, 2nd respectivelw n-th compenent expenenttal variobles with mean Let X denete the length of time until faihune of the sustem The cempsnents are comected in sentes - The sustaxn feuls when at least one ef the n cempenents Foils .-. Length ef time wntil the system fails = Lengti et time until at least one component fails Thus accordm to the problem it can be salà thob × X2 lf 2nd compon ent fails Art x Xn n-th cemponent foils £rst
乂/.서 Variance of X =E(X)-EON し(x*)= 분 еーカx/u.du . 2バ (カ 2

2c)

x denotes the measur ement of eye pressure Dn a Onofes Person 다 a person has guau corn a 다 α Person does not have glaucoma I tccordinao he problem:(i)

Requived probabilb - (x-25)72. 2N 2下 2丌 2.(ii)